1 15 oints) Deterine if the following propositions are TRUE or FALSE. Note that p, q r are propositi Px) and P(x.y) are predicates. RUE or FALSE.Note that p, q, r are propositions. (a).TNE 1f2小5or I + 1-3, then 10+2-3or 2 + 2-4. (b).TRvE+1 0 if and only if 2+ 2 5. (d). _ p v T Ξ T, where p is a proposition and T is tautology. V x Px) is equivalent to Vx - Px) (g). ㅡㅡㅡ, y P(x, y) is equivalent to Vy Vx P(x,y). w x Vx 3 y P(x, y) is equivalent to 3 y Vx P(x.y). I +2=0,XOR2 + 2 = 4 p ^ q v F Ξ F, where F is contradiction. A predicate is a propositional function. (k),
Answer #1

(c) ("5 > 6" intersection "1 + 1 = 3") union "7 > 5"

(5 > 6 - (false ) intersection 1 + = 3 (false) ) unon 7 > 5 (true) => (false intersection false) union (true) => Hence answer is true

(i) 1 + 2 = 0 XOR 2 + 2 = 4 - False, In XOR 2 + 2 = 0

(l) a predicate is propositional function - True, because from the predicate or propositional forma the propositional functions has been extracted.